OpenStudy (anonymous):

PLEASE PLEASE PLEASE PLEASE HELP!!! How can I solve an exponetial model to make an inverse one ?

OpenStudy (tkhunny):

That would depend on the exponential model. Please provide one and give it your best effort.

OpenStudy (anonymous):

h(t)=106(0.95)^t+50

OpenStudy (anonymous):

I got as far as \[t-50\div160=0.95^{h}\]

OpenStudy (tkhunny):

Pretty good, excepting that you have forgotten your Order of Operations. I hope you mean \((t-50)/160 = 0.95^{h}\). Without the parentheses, it is quite a different matter. Have you considered introducing a logarithm?

OpenStudy (anonymous):

Yes, which is where I get confused

OpenStudy (anonymous):

log\[\log_{0.95}(t-50)\div160=\log_{0.95}(0.95)^h\]

OpenStudy (tkhunny):

Logs base 0.95?!! Why on earth would you do that? Use somethin normal. \(\log\left(\dfrac{t-50}{160}\right) = \log\left(0.95^{h}\right)\) Now, use logarithm rules. \(\log\left(\dfrac{t-50}{160}\right) = h\cdot\log\left(0.95\right)\) You're almost done. Note: You can use the logs base 0.95 if you want, but I'm guessing that will be more confusing.

OpenStudy (anonymous):

Okay so for the the side with H I would get -0.022

OpenStudy (anonymous):

How do I get rid of that ?

OpenStudy (anonymous):

?

OpenStudy (tkhunny):

Why did you get that at all? \(\log(0.95)\) is just fine.

OpenStudy (anonymous):

and that would be my inverse function ?

OpenStudy (tkhunny):

Well, did you solve for \(h\)?

OpenStudy (anonymous):

h=1/160 log(t-50)/log(.95)

OpenStudy (anonymous):

Is that simplified enough ?

OpenStudy (tkhunny):

?? How did the 160 get outside the logarithm? If you were to use logarithm rules, you should get \(\log((t-50)/160) = \log(t-50) - log(160)\).

OpenStudy (tkhunny):

You can also use the base transformation to rewrite a little, but it's still not very pretty. \(\dfrac{\log\left(\dfrac{t-50}{160}\right)}{\log(0.95)} = \log_{0.95}\left(\dfrac{t-50}{160}\right) = h\)

OpenStudy (anonymous):

okay thank you